A solution of 60 ml is 50% acid. How many ml of water must be added to create a 30% acid solution?
I tried solving this problem but got a negative answer...
I put the question in this form- W+.50(60)=.30(W+60)
Thank you.

Answer by Alan3354(69443)

A solution of 60 ml is 50% acid. How many ml of water must be added to create a 30% acid solution?
-----------
60 ml of 50% has 30 ml of acid.
For 30 ml to be 30%, the total must be 100 ml
Add 40 ml of water.
--------------------
You can do it this way:
T = w + 60 (total solution)
0.5*60 + w*0 = 0.3(w + 60) (total acid)
30 = 0.3w + 18
0.3w = 12
w = 40 ml